OpenStudy (anonymous):

the function f(x)=9x+8/x-1 is one to one, find an equation for f^-1(x)

OpenStudy (anonymous):

let x=9y+8/y-1 x(y-10=9y+8 xy-x=9y+8 xy-9y=8+x y(x-9)=8+x/x-9

OpenStudy (anonymous):

thank you!!